OpenStudy (anonymous):

write the slope intercept quation for the line that passes through (-4, 8) and (-3, -2)

OpenStudy (saifoo.khan):

y = -10x-32

OpenStudy (anonymous):

^

OpenStudy (anonymous):

saifoo.khan: how did you come to that answer?

OpenStudy (anonymous):

(-4, 8) and (-3, -2) m = (y2 - y1)/(x2 - x1) = (-2-8)/(-3+4) = -10 = slope y = mx + b to find b sub one of the given co-ordinates, and the slope -2 = (-10)(-3) + b -2 = 30 + b b= -32 :- y = -10x -32